Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 17     Entry time: 09 Feb 2024, 07:15
Author: Giuseppe Mercurio 
Group: SCS 
Subject: Shift summary 

night shift summary

achievements:

  • characterized FEL beam size to have 20 um x 20 um (Hor x Ver)
  • exchanged sample
  • closed KB baffle, HSLITS, guardstage
  • started fluence scan using OL on the membranes

problems:

  • sometimes the macro does not react when we press 'experiment', after shutting down and restarting the macro, it works

plan:

  • continue with scans on the membranes.
ELOG V3.1.4-7c3fd00